How many grams of lithium hydride are necessary to completely react with 10.3 grams of titanium(IV) chloride in the reaction below?

TiCl4(l) + 4LiH(s) Ti(s) + 4LiCl(s) + 2H2(g)
1
Expert's answer
2020-03-31T13:42:50-0400

m(LiH)=m(TiCl4)*M(LiH)*4/M(TiCl4)=10.3(g)*7.95(g/mol)*4/189.68(g/mol)=1.73g